A standard 100 W (120 V) lightbulb contains a 9.10 cm -long tungsten filament. The high-temperature resistivity of tungsten is 9.0×10−7Ωm. What is the diameter of the filament?
Answer #1

We know power of a bulb

P=\frac{V^2}{R}

where V is voltage and R is the resistance of the filament.

Here P =100 watt, V=120 volt ,so

R=\frac{V^2}{P}=\frac{(120)^2}{100}=144 ohm

Again for a filament of D diameter, l length and high-temperature resistivity \rho, the resistance of the filament is

R=\rho\cdot\frac{l}{A}=\rho\cdot\frac{l}{\pi (\frac{D}{2})^2}

\Rightarrow D=2\times\sqrt{\frac{\rho l}{\pi R}}

Here l=9\cdot10 cm, l=9\cdot10\times10^-2 meter ,\rho=9\cdot0\times10^-7\Omega-m

So D=2\times\sqrt{\frac{9\cdot0\times10^-7\times9\cdot10\times10^-2}{\pi\times144}}=2\cdot69\times10^-5 meter